# Conformation and dynamics of wet tangentially-driven active filaments

### v2

@@ Abstract
We explore the impact of hydrodynamic interactions on the conformational and dynamical properties of wet tangentially-driven active polymers using multiparticle collision dynamics simulations. By analyzing active filaments with varying degrees of flexibility, we find that fluid-mediated interactions significantly influence both their conformation and dynamics. These interactions cause polymer conformations to shrink, especially for semiflexible polymers at high activity levels, where the average size of wet chains becomes nearly three times smaller, due to formation of helix-like structures. This hydrodynamic-induced shrinkage is a hallmark of active polymers, as fluid-mediated interactions have a minimal effect on the mean conformation of passive polymers. Furthermore, for tangentially-driven polymers where activity and conformation are coupled, hydrodynamic interactions significantly enhance the orientational and translational dynamics compared to their dry counterparts.
@@ I Introduction
Active filaments and polymers can be found on a wide range of scales from ATP-powered chromatin in cell nucleus and motor-driven actin and microtubules to microscopic worms and snakes. Inspired by experiments of active biological filaments a diverse range of theoretical and computational models for active polymers have been proposed in recent years Winkler et al. (2017); Winkler and Gompper (2020) including scalar Smrek and Kremer (2018); Chubak et al. (2020) and force-driven active polymer models where the active force can be stochastic Eisenstecken et al. (2016); Ghosh and Gov (2014), tangential Isele-Holder et al. (2015); Bianco et al. (2018) or perpendicular Prathyusha et al. (2022) to the backbone. Among these, the tangentially-driven active polymer model has garnered increased interest due to its relevance to filament-like systems, such as motor-driven actin and microtubule filaments, elongated bacteria like Proteus mirabilis Lough et al. (2023), and worm species such as T. tubifex  Deblais et al. (2020) that propel themselves along their backbone.
@@
The majority of studies on active polymers have ignored hydrodynamic interactions and have focused on the interplay between internal dynamics and activity mechanisms. There are a few studies that investigate the role of hydrodynamic interactions on the structural and dynamical properties of active filaments Zöttl and Yeomans (2019); Jayaraman et al. (2012); Laskar and Adhikari (2015); De Canio et al. (2017); Saintillan et al. (2018); Martín-Gómez et al. (2019, 2019); Mahajan and Saintillan (2022); Stein et al. (2021); Lough et al. (2023). These include studies of active Brownian polymers in solution Martín-Gómez et al. (2019, 2019) and active polymer models where monomers are subject to permanent or stochastic active force dipoles along their backbone Jayaraman et al. (2012); Mahajan and Saintillan (2022). These investigations reveal the significance of hydrodynamic interactions on polymer dynamics and even on conformational properties.
@@
In this work, we analyze the conformational and dynamical properties of tangentially-driven active polymers Isele-Holder et al. (2015) embedded in a fluid where the monomers are externally actuated, focusing on unconfined dilute active polymer solutions. The dry version of this model has been subject of intensive studies Isele-Holder et al. (2015); Fazelzadeh et al. (2023); Bianco et al. (2018); Mokhtari and Zippelius (2019); Karan et al. (2024); Khalilian et al. (2024); Duman et al. (2018); Prathyusha et al. (2018); Vatin et al. (2024). This model is relevant for biological filaments actuated by molecular motors and chains of Janus active colloids bonded by electrostatic or magnetic dipolar interactions Yan et al. (2016); Nariaki and C. Patrick (2020); Chao et al. (2024) resulting in an active force tangential to the backbone. Comparing the conformation and dynamics of dry and wet tangentially-driven polymers we address two important questions. (i) Do the hydrodynamic interactions lead to the emergence of new modes of conformation and transport of active polymers? (ii) How do the longtime dynamics of the wet active polymers differ from that of the dry ones?
@@
To account for hydrodynamic interactions, we employ the explicit fluid simulation method of multi particle collision dynamics (MPCD). The MPCD is a mesoscale particle-based simulation method for simulation of fluids, which naturally incorporates thermal fluctuations. The MPCD is able to capture the hydrodynamic correlations correctly Malevanets and Kapral (1999); Howard et al. (2019); Tao et al. (2008); Llahí et al. (2022); Huang et al. (2012) and is widely used for simulating dynamics of colloidal and polymeric solutions. The MPCD method allows to faithfully reproduce the fluid-mediated interactions with significantly less computational costs. In the context of active systems, it has been utilized to study the dynamics of microswimmers such as generic squirmer model Babu and Stark (2012); Hu et al. (2015); Downton and Stark (2009); Götze and Gompper (2010); Zöttl and Stark (2018) and wet active Brownian polymer models Martín-Gómez et al. (2019, 2019). In this article, we present an implementation of wet tangentially-driven polymers in MPCD. The polymer model consists of linear bead-spring chains where each monomer experiences an external active force tangential to the local tangent of the backbone Isele-Holder et al. (2015).
@@
We investigate the conformational and dynamical features of wet tangentially-driven polymers with varying activity levels and two different degrees of flexibility, comparing them with those of their dry counterparts. Our investigations reveal that fluid-mediated interactions influence both the conformational and dynamical properties remarkably. Hydrodynamic interactions cause polymer conformations to shrink, an effect that becomes more pronounced with increasing activity levels and bending stiffness. At high activity levels, the conformations of semiflexible polymers becomes helix-like due to consecutive extensional and compressional flow profiles created by active forces around the polymer backbone which compete with its bending elasticity. Wet active polymers exhibit faster orientational dynamics and enhanced longtime diffusion coefficient. Interestingly, the orientational relaxation time of the end-to-end vector of the polymers is almost independent of their degree of flexibility but decreases with increasing active force.
@@ II.1 Multiparticle Collision Dynamics simulations
In the MPC approach, the fluid is simulated by many identical point-like particles, which undergo discrete iterative streaming and collision steps. During each step, the MPC particles propagate and locally interact with each other Malevanets and Kapral (1999). In the streaming step the fluid particles of mass $m$ are propagated by solving Newton’s equations of motion over a small time step $\delta t$ using the Velocity-Verlet integration scheme given by
$\begin{split}\mathbf{\bm{v}}_{i}(t+\frac{1}{2}\delta t)=\mathbf{\bm{v}}^{\prime}_{i}(t)+\frac{\mathbf{\bm{f}}_{i}}{2m}\delta t\\ \mathbf{\bm{r}}_{i}(t+\delta t)=\mathbf{\bm{r}}_{i}+\delta t\mathbf{\bm{v}}_{i}(t+\frac{1}{2}\delta t)\\ \mathbf{\bm{v}}_{i}(t+\delta t)=\mathbf{\bm{v}}_{i}(t+\frac{1}{2}\delta t)+\frac{\mathbf{\bm{f}}}{2m}\delta t\end{split}$ (1)
Here, $v_{i}$ and $r_{i}$ are the velocity and position of monomer $i$ respectively, $v^{\prime}_{i}$ is its updated velocity vector after multiparticle collision step given by Eq. (2), and $\mathbf{\bm{f}}_{i}$ is the force acting on the particle $i$. After streaming, the solvent undergoes a multiparticle collision representing the core of the MPC algorithm. To implement the collision steps the fluid particles are sorted into cubic cells of edge length $L_{cell}$, which sets the spatial resolution of HI. Thereafter, the relative velocity of each particle with respect to the centre of mass velocity of all the particles in the cell is rotated by some constant angle $\alpha$ about a randomly selected axis. This collision method is called Stochastic Rotational Dynamics (SRD) Ihle and Kroll (2001). Update of velocities of all particles in the same cell undergoing a stochastic collision is described by
@@
where $\mathbf{\mathcal{R}}(\alpha)$ is a rotation matrix and $\mathbf{\bm{v}}_{cm}=\frac{1}{N_{c}}\sum^{N_{c}}_{i=1}\mathbf{\bm{v}}_{i}$ is the centre of mass velocity of a collision cell. The updated velocities $\mathbf{\bm{v}}^{\prime}_{i}(t+\delta t)$ are then used for repeating the streaming step described by Eq. 1 .
Figure 1: Schematics of the active tangentially driven polymer, MPC particles and MPC mesh. The polar nature of this particular type of activity let us define a head and a tail for the filament.
@@ II.2 Active polymer model
is the sum of the intramolecular potentials acting on the $i$th monomer. Likewise, $\mathbf{\bm{F}}^{a}_{i}$ and $\mathbf{\bm{F}}^{H}_{i}$ describe the active force and hydrodynamic force acting on the monomer. The latter force results in from the interaction of solvent molecules with the monomers and its effect will be incorporated by coupling the dynamics of monomers to the MPC fluid particles as will be discussed in the subsequent section.
@@
where $f^{a}$ sets the strength of the active force. The active force on the tail monomer is given by $\mathbf{\bm{F}}^{a}_{1}=\frac{f^{a}}{2\ell}\mathbf{\bm{R}}_{1,2}$ and for the head monomer is $\mathbf{\bm{F}}^{a}_{N_{m}}=\frac{f^{a}}{2\ell}\mathbf{\bm{R}}_{N_{m}-1,N_{m}}$. To incorporate the hydrodynamic force on each monomer, we couple it to the MPC particles as discussed in the following section.
@@ II.3 Coupling of polymer dynamics to fluid
where $\mathbf{\bm{F}}^{a}_{tot}$ is the sum of all the tangential forces applied to the monomers. The sum of forces induced by intramolecular interactions in the polymer cancel out due to Newton’s third law, thus we can focus purely on the total active force $\mathbf{\bm{F}}^{a}_{tot}$ itself. In confined systems, the global fluid flow creates a backflow. We can achieve a similar backflow in systems with periodic boundary conditions by applying an additional backflow force to all monomers and MPC particles so that the total force becomes null, i.e. $\mathbf{\bm{F}}^{a}_{tot}(t)+\mathbf{\bm{F}}^{b}_{tot}(t)=0$. We distribute the backflow forces acting on the monomers $\mathbf{\bm{F}}^{b}_{mon}$ and MPC particles $\mathbf{\bm{F}}^{b}_{MPC}$ according to their masses Martin-Gomez et al. (2020); Llahí et al. (2022) as
@@
Here, $M$ and $m$ are the masses of the monomers and MPC particles and $N_{m}$ and $N$ denote the number of monomers and MPC particles respectively. The magnitude of the backflow force is governed by the ratio between the particle mass and the total mass of the particles in the system. Assuming that the timescale of changes of polymer conformation are larger than the collision time $\delta t$, the MPC fluid particle velocities $\mathbf{\bm{v}}_{k}$ and positions $\mathbf{\bm{r}}_{k}$ after a streaming step are given by
@@
which incorporates coupling of monomers dynamics to the background flow. Here, $\mathbf{\bm{F}}_{mon}^{b}$ is the backflow force applied to each monomer given by Eq. (10). Eq. 13 is solved using the velocity Verlet integration scheme similar to Eq. (12) but using the total force acting on each monomer instead of $F^{b}_{MPC}$. The SRD collision step now requires the fluid particles and monomers to undergo stochastic rotations relative to the centre of mass velocity of the collision cells which is given by
@@
where $N_{c}$ and $N^{m}_{c}$ are the number of fluid particles and monomers in a collision cell respectively. The stochastic rotation of MPC fluid particles and monomers take place according to
@@
where $\mathbf{\bm{v}}_{k}(t+\delta t)$ is determined by Eq. (12) and $\mathbf{\bm{V}}_{i}(t+\delta t)$ is given by the solution of Eq. (13) obtained by the velocity Verlet algorithm. This implies that a Stokeslet flow field by the active force appears for every monomer. The discussed algorithm for external actuated active polymers with tangential active force is implemented in Hoomd-blue Glaser et al. (2015); Anderson et al. (2008) with a home-made modification to include the active and the backflow forces for the coupling of the MPC fluid to the monomer dynamics.

### v3

@@ Abstract
We explore the impact of hydrodynamic interactions on the conformational and dynamical properties of wet tangentially-driven active polymers using multiparticle collision dynamics simulations. By analyzing active filaments with varying degrees of flexibility, we find that fluid-mediated interactions significantly influence both their conformation and dynamics. These interactions cause polymer conformations to shrink relative to their dry counterparts, especially for semiflexible polymers at high activity levels, where the average size of wet chains becomes nearly three times smaller, due to local buckling of wet polymers. This hydrodynamic-induced shrinkage is a hallmark of active polymers, as fluid-mediated interactions do not affect conformational properties of passive polymers. Furthermore, for tangentially-driven polymers where activity and conformation are coupled, hydrodynamic interactions significantly enhance the orientational and translational dynamics compared to their dry counterparts.
@@ I Introduction
Active filaments and polymers can be found on a wide range of scales from ATP-powered chromatin in cell nucleus and motor-driven actin and microtubules to microscopic worms and snakes. Inspired by experiments of active biological filaments a diverse range of theoretical and computational models for active polymers have been proposed in recent years Winkler et al. (2017); Winkler and Gompper (2020) including scalar Smrek and Kremer (2018); Chubak et al. (2020) and force-driven active polymer models where the active force can be stochastic Eisenstecken et al. (2016); Ghosh and Gov (2014), tangential Isele-Holder et al. (2015); Bianco et al. (2018); Dreyfus et al. (2005); Tejedor et al. (2024) or perpendicular Prathyusha et al. (2022) to the backbone. Among these, the tangentially-driven active polymer model has garnered increased interest due to its relevance to filament-like systems, such as motor-driven actin and microtubule filaments, elongated bacteria like Proteus mirabilis Lough et al. (2023), and worm species such as T. tubifex  Deblais et al. (2020) that propel themselves along their backbone.
@@
The majority of studies on active polymers have ignored hydrodynamic interactions and have focused on the interplay between internal dynamics and activity mechanisms. There are a few studies that investigate the role of hydrodynamic interactions on the structural and dynamical properties of active filaments Lough et al. (2023); Barriuso Gutiérrez et al. (2022); Balin et al. (2017); Zöttl and Yeomans (2019); Jayaraman et al. (2012); Laskar and Adhikari (2015); De Canio et al. (2017); Saintillan et al. (2018); Martín-Gómez et al. (2019, 2019); Mahajan and Saintillan (2022); Stein et al. (2021); Jiang and Hou (2014). These include studies of active Brownian polymers in solution Martín-Gómez et al. (2019, 2019) and active polymer models where monomers are subject to permanent or stochastic active force dipoles along their backbone Jayaraman et al. (2012); Mahajan and Saintillan (2022). These investigations reveal the significance of hydrodynamic interactions on polymer dynamics and even on conformational properties.
@@
In this work, we analyze the conformational and dynamical properties of tangentially-driven active polymers Isele-Holder et al. (2015) embedded in a fluid where the monomers are externally actuated, focusing on unconfined dilute active polymer solutions. The dry version of this model has been subject of intensive studies Isele-Holder et al. (2015); Fazelzadeh et al. (2023); Bianco et al. (2018); Mokhtari and Zippelius (2019); Karan et al. (2024); Khalilian et al. (2024); Duman et al. (2018); Prathyusha et al. (2018); Vatin et al. (2024).
Biological active filaments driven by molecular motors, as well as chains of Janus active colloids bonded through electrostatic or magnetic dipolar interactions Yan et al. (2016); Nariaki and C. Patrick (2020); Chao et al. (2024), are examples of systems where the active force is constrained to be tangential to the backbone. These systems are force-neutral and, at the hydrodynamic level, can be described as chains of hydrodynamic dipoles aligned tangentially to the backbone, similar to the models developed in Jayaraman et al. (2012); Mahajan and Saintillan (2022). In contrast, active polymers in motility assays driven by molecular motors fixed to the assay are not force-neutral and they can be considered as externally actuated Kawamura et al. (2008); Keya et al. (2020). Moreover, externally-actuated active polymers with tangential driving can be engineered by applying a locally tangential drive to a chain of colloids with flexible linkers using dynamic optical tweezeing Curtis et al. (2002); Eriksen et al. (2002); Ghosh and Ghosh (2019).
Comparing the conformation and dynamics of dry and wet tangentially-driven polymers we address two important questions. (i) Do the hydrodynamic interactions lead to the emergence of new modes of conformation and transport of active polymers? (ii) How do the longtime dynamics of the wet active polymers differ from that of the dry ones?
@@
To account for hydrodynamic interactions, we employ the explicit fluid simulation method of multi particle collision dynamics (MPCD). The MPCD is a mesoscale particle-based simulation method for simulation of fluids, which naturally incorporates thermal fluctuations. The MPCD is able to capture the hydrodynamic correlations correctly Malevanets and Kapral (1999); Howard et al. (2019); Tao et al. (2008); Clopés Llahí et al. (2022); Huang et al. (2012) and is widely used for simulating dynamics of colloidal and polymeric solutions Slater et al. (2009); Malevanets and Kapral (1999); Shendruk and Yeomans (2015); Howard et al. (2018); Lamura et al. (2001); Lamura (2022); Bolintineanu et al. (2014); Ripoll et al. (2007); Ruiz-Franco et al. (2019); Wani et al. (2022); Hecht et al. (2005); Padding and Louis (2006); Gompper et al. (2009). The MPCD method allows to faithfully reproduce the fluid-mediated interactions with significantly less computational costs. In the context of active systems, it has been utilized to study the dynamics of microswimmers such as generic squirmer model Babu and Stark (2012); Hu et al. (2015); Downton and Stark (2009); Götze and Gompper (2010); Zöttl and Stark (2018) and wet active Brownian polymer models Martín-Gómez et al. (2019, 2019). In this article, we present an implementation of wet tangentially-driven polymers in MPCD. The polymer model consists of linear bead-spring chains where each monomer experiences an external active force tangential to the local tangent of the backbone Isele-Holder et al. (2015).
@@
We investigate the conformational and dynamical features of wet tangentially-driven polymers with varying activity levels and two different degrees of flexibility, comparing them with those of their dry counterparts. Our investigations reveal that fluid-mediated interactions influence both the conformational and dynamical properties remarkably. Hydrodynamic interactions cause polymer conformations to shrink, an effect that becomes more pronounced with increasing activity levels and bending stiffness. At high activity levels, the conformations of semiflexible polymers buckle and locally fold due to consecutive extensional and compressional flow profiles created by active forces around the polymer backbone which compete with its bending elasticity. The conformations of wet semiflexible active polymers resemble qualitatively to buckled conformations of microtubules in motility assays Kawamura et al. (2008); Keya et al. (2020), which form rings and spirals. Wet active polymers exhibit faster orientational dynamics and enhanced longtime diffusion coefficient. Interestingly, the orientational relaxation time of the end-to-end vector of the polymers is almost independent of their degree of flexibility but decreases with increasing active force.
@@ II.1 Multiparticle Collision Dynamics simulations
In the MCPD approach, the fluid is simulated by many identical point-like particles, which undergo discrete iterative streaming and collision steps. During each step, the MCPD particles propagate and locally interact with each other Malevanets and Kapral (1999). In the streaming step the fluid particles of mass $m$ are propagated by solving Newton’s equations of motion over a small time step $\delta t$ using the Velocity-Verlet integration scheme given by
$\begin{split}\mathbf{\bm{v}}_{i}(t+\frac{1}{2}\delta t)=\mathbf{\bm{v}}^{\prime}_{i}(t)+\frac{\mathbf{\bm{f}}_{i}}{2m}\delta t\\ \mathbf{\bm{r}}_{i}(t+\delta t)=\mathbf{\bm{r}}_{i}+\delta t\mathbf{\bm{v}}_{i}(t+\frac{1}{2}\delta t)\\ \mathbf{\bm{v}}_{i}(t+\delta t)=\mathbf{\bm{v}}_{i}(t+\frac{1}{2}\delta t)+\frac{\mathbf{\bm{f}}_{i}}{2m}\delta t\end{split}$ (1)
Here, $v_{i}$ and $r_{i}$ are the velocity and position of monomer $i$ respectively, $v^{\prime}_{i}$ is its updated velocity vector after multiparticle collision step given by Eq. (2), and $\mathbf{\bm{f}}_{i}$ is the force acting on the particle $i$. After streaming, the solvent undergoes a multiparticle collision representing the core of the MCPD algorithm. To implement the collision steps the fluid particles are sorted into cubic cells of edge length $L_{cell}$, which sets the spatial resolution of HI. Thereafter, the relative velocity of each particle with respect to the centre of mass velocity of all the particles in the cell is rotated by some constant angle $\alpha$ about a randomly selected axis. This collision method is called Stochastic Rotational Dynamics (SRD) Malevanets and Kapral (1999); Ihle and Kroll (2001). The choice of collision rule and its parameters determines the transport coefficients of the mesoscale MPCD fluid Kikuchi et al. (2003). Update of velocities of all particles in the same cell with $N_{c}$ MPCD particles undergoing a stochastic collision is described by
@@
where $\mathbf{\mathcal{R}}(\alpha)$ is a stochastic rotation matrix with angle $\alpha$ and $\mathbf{\bm{v}}_{cm}=\frac{1}{N_{c}}\sum^{N_{c}}_{i=1}\mathbf{\bm{v}}_{i}$ is the centre of mass velocity of a collision cell. The updated velocities $\mathbf{\bm{v}}^{\prime}_{i}(t+\delta t)$ are then used for repeating the streaming step described by Eq. 1. The SRD method is not isothermal on its own as it naturally imparts the NVE ensemble to the system. To apply a thermostat with constant temperature $T$, we use a Maxwell-Boltzmann-scaling (MBS) approach Huang et al. (2015). This thermostat rescales the relative velocities $\Delta\mathbf{\bm{v}}_{i}=\mathbf{\bm{v}}_{i}-\mathbf{\bm{v}}_{cm}$ of the particles in a collision cell by a constant factor $\mu$ after each collision step using Maxwell-Boltzmann distribution of kinetic energy. The rescaling factor $\mu$ typically differs for every cell and collision-time step. Since the total relative momentum of a collision cell is zero, such a scaling leaves the total momentum of a cell intact. The MBS has been proven to render the correct hydrodynamics properties of isothermal fluids Huang et al. (2010). To obtain $\mu$ for each collision cell, at every collision, a random kinetic energy $E_{k}^{{}^{\prime}}$ is taken from the distribution function of kinetic energy of the MPCD ideal-gas particles given by Maxwell-Boltzmann distribution Huang et al. (2010, 2015):
$P(E_{k}^{{}^{\prime}})=\frac{1}{E_{k}^{{}^{\prime}}\Gamma(\frac{f}{2})}\Bigg{(}\frac{E_{k}^{{}^{\prime}}}{k_{B}T}\Bigg{)}^{\frac{f}{2}}\exp\Bigg{(}-\frac{E_{k}^{{}^{\prime}}}{k_{B}T}\Bigg{)}.$ (3)
Here, $\Gamma$ denotes the Gamma function and $f=3N_{c}-1$ is the number of the degrees of freedom of the cell and the mean kinetic energy is given by $\langle E_{k}\rangle=f/2k_{B}T$.
Then, the respective scale factor for the velocities is obtained as
$\mu=\Bigg{(}\frac{E_{k}^{{}^{\prime}}}{E_{k}}\Bigg{)}^{\frac{1}{2}},$ (4)
where $E_{k}=m/2\sum_{i=1}^{N_{c}}\Delta\mathbf{\bm{v}}_{i}^{2}$ is the kinetic energy in the center of mass frame of each cell before rescaling.
Figure 1: Schematics of the active tangentially driven polymer,MCPD particles andMCPD mesh. The polar nature of this particular type of activity let us define a head and a tail for the filament.
@@ II.2 Active polymer model
is the sum of the intramolecular potentials acting on the $i$th monomer. Likewise, $\mathbf{\bm{F}}^{a}_{i}$ and $\mathbf{\bm{F}}^{H}_{i}$ describe the active force and hydrodynamic force acting on the monomer. The latter force results in from the interaction of solvent molecules with the monomers and its effect will be incorporated by coupling the dynamics of monomers to theMCPD fluid particles as will be discussed in the subsequent section.
@@
where $f^{a}$ sets the strength of the active force. The active force on the tail monomer is given by $\mathbf{\bm{F}}^{a}_{1}=\frac{f^{a}}{2\ell}\mathbf{\bm{R}}_{1,2}$ and for the head monomer is $\mathbf{\bm{F}}^{a}_{N_{m}}=\frac{f^{a}}{2\ell}\mathbf{\bm{R}}_{N_{m}-1,N_{m}}$. To incorporate the hydrodynamic force on each monomer, we couple it to the MCPD particles as discussed in the following section.
@@ II.3 Coupling of polymer dynamics to fluid
where $\mathbf{\bm{F}}^{a}_{tot}$ is the sum of all the tangential forces applied to the monomers. The sum of forces induced by intramolecular interactions in the polymer cancel out due to Newton’s third law, thus we can focus purely on the total active force $\mathbf{\bm{F}}^{a}_{tot}$ itself.
For an externally-driven active polymer in real experimental situations with confining wall, the net active force $\mathbf{\bm{F}}^{a}_{tot}$ exerted on the fluid is reflected by the confining container walls and induces a backflow. The whole system has zero momentum because the walls do not let the external force to accelerate the center of mass of the whole system. In contrast, when we impose periodic boundary conditions in simulation, MCPD fluid particles can accelerate due to the external active force and result in a net fluid flow which accelerates the whole system in the absence of physical walls. A globally moving system is an artifact of the periodic boundary conditions and not physical. To prevent a net fluid flow and artificial acceleration of center of mass of the whole system, we apply a decelerating backflow force to all the monomer and MPCD particles in such a way that the total momentum of the system (fluid plus active polymer) remains zero, i.e., $\mathbf{\bm{F}}^{a}_{tot}(t)+\mathbf{\bm{F}}^{b}_{tot}(t)=0$. We distribute backflow force equally among all MPCD particles and monomers proportional to their mass to produce a uniform acceleration opposite to the direction of total active force, similar to the case of sedimenting soft colloids exposed to a uniform gravitational force Singh et al. (2018) or externally driven active Brownian polymers Martin-Gomez et al. (2020); Clopés Llahí et al. (2022).
More explicitly the distribution of backflow forces acting on the monomers $\mathbf{\bm{F}}^{b}_{mon}$ and MCPD particles $\mathbf{\bm{F}}^{b}_{MPC}$ according to their masses Martin-Gomez et al. (2020); Clopés Llahí et al. (2022) are given by
@@
Here, $M$ and $m$ are the masses of the monomers and MCPD particles and $N_{m}$ and $N$ denote the number of monomers and MCPD particles respectively. The magnitude of the backflow force is governed by the ratio between the particle mass and the total mass of the particles in the system.
Assuming that the timescale of changes of polymer conformation are larger than the collision time $\delta t$, the MCPD fluid particle velocities $\mathbf{\bm{v}}_{k}$ and positions $\mathbf{\bm{r}}_{k}$ after a streaming step are given by
@@
which incorporates coupling of monomers dynamics to the background flow. Here, $\mathbf{\bm{F}}_{mon}^{b}$ is the backflow force applied to each monomer given by Eq. (12). Eq. 15 is solved using the velocity Verlet integration scheme similar to Eq. (14) but using the total force acting on each monomer instead of $F^{b}_{MPC}$. The SRD collision step now requires the fluid particles and monomers to undergo stochastic rotations relative to the centre of mass velocity of the collision cells. The stochastic collisions including both MPCD particles and the monomers account for interactions of solute with solvent and a build up of hydrodynamic interactions.
The center of mass of which is given by
@@
where $N^{m}_{c}$ is the number of monomers in a collision cell. The stochastic rotation of MCPD fluid particles and monomers take place according to
@@
where $\mathbf{\bm{v}}_{k}(t+\delta t)$ is determined by Eq. (14) and $\mathbf{\bm{V}}_{i}(t+\delta t)$ is given by the solution of Eq. (15) obtained by the velocity Verlet algorithm. This implies that a Stokeslet flow field by the active force appears for every monomer.
The discussed algorithm for external actuated active polymers with tangential active force is implemented in Hoomd-blue Glaser et al. (2015); Anderson et al. (2008) with a home-made modification to include the active and the backflow forces for the coupling of the MCPD fluid to the monomer dynamics.